OpenStudy (anonymous):

Now Lets So Who Can Answer This ! Use the distributive property to factor the expression below. 3x - 9x2

OpenStudy (anonymous):

3x( 1-3x) ?

OpenStudy (ash2326):

We have \[3x-9x^2\] According to distributive property \[A(B+C)=AB+AC\] so here we have \[3x\times 1- 3x\times 3x\] Now using the reverse of distributive property \[ 3x(1-3x)\] And voila we have our factors!!!